drug concentration: 40 mg/ml, 20ml per vial

dose: 70mg/kg/24hr

patient weight: 65 kg

1. What is the rate of infusion(ml/hr) that should be entered on the pharmacy label?

1a How many vials of drug would be used for the patient? please show work
Answer #1

ANSWER

1. Rate of infusion (ml / hr)

    * Formula

           Rate of infusion = Volume (ml) / Time (hr)

      Ordered dose = 70mg/kg/24hr.

          Patient body weight = 65 kg.

           Total volume needed for the patient = 70 * 65

                                                                    = 4550 mg.

                                                          1 ml = 40 mg.(drug concentration)

                                  So Volume = 4550 / 40        

                                                                   = 113.8 ml.

                                                   Total time = 24 hr.

                                         Rate of infusion = 113.8 / 24

                                                                     = 4.74 ml/ hr.

* Number of vials of drug used for the patient

                                      Drug concentration = 40mg /ml.

                                                             1 vial = 20 ml.

             Total volume needed for the patient = 113.8 ml

           Total vials needed for the patient = 113.8 / 20

                                                                       = 5.7

                                                                      = 6 vials.